# Traversability-Aware Legged Navigation by Learning from Real-World Visual Data

### v1

@@ Abstract
The enhanced mobility brought by legged locomotion empowers quadrupedal robots to navigate through complex and unstructured environments. However, optimizing agile locomotion while accounting for the varying energy costs of traversing different terrains remains an open challenge. Most previous work focuses on planning trajectories with traversability cost estimation based on human-labeled environmental features. However, this human-centric approach is insufficient because it does not account for the varying capabilities of the robot locomotion controllers over challenging terrains. To address this, we develop a novel traversability estimator in a robot-centric manner, based on the value function of the robot’s locomotion controller. This estimator is integrated into a new learning-based RGBD navigation framework. The framework develops a planner that guides the robot in avoiding obstacles and hard-to-traverse terrains while reaching its goals. The training of the navigation planner is directly performed in the real world using a sample efficient reinforcement learning method. Through extensive benchmarking, we demonstrate that the proposed framework achieves the best performance in accurate traversability cost estimation and efficient learning from multi-modal data (the robot’s color and depth vision, and proprioceptive feedback) for real-world training. Using the proposed method, a quadrupedal robot learns to perform traversability-aware navigation through trial and error in various real-world environments with challenging terrains that are difficult to classify using depth vision alone.
@@ I Introduction
Figure 1: Our proposed framework enables a quadrupedal robot to learn to avoid hard-to-traverse terrains (such as muddy areas covered by leaves) and obstacles from its own interactions in the real world. The robot learns to utilize its onboard color and depth vision sensors to identify challenging terrains and obstacles while navigating toward the goals, marked by the blue dashed circles. The inclusion of RGB images allows the robot to identify additional terrain textures that are difficult to perceive with depth images alone during the navigation.
However, estimating the traversability cost of terrain and planning to avoid high-cost regions for legged robots is a challenging problem. It requires the robot to interpret visual textures and corresponding physical properties, such as friction and elasticity, of different terrains. For example, when humans encounter a muddy road with high traversability costs, we first see the mud’s texture and infer its difficulty based on previous experience. We highlight the importance of learning from previous experience as it’s difficult to accurately estimate the terrain’s actual cost until we traverse it ourselves. For legged robots to develop similar intelligence, the robot needs to fuse color data (terrain texture), depth data (identifying non-traversable regions like obstacles), and proprioceptive feedback (implicitly encoding the physics) in real time and the real world. This problem is further complicated by the complex dynamics of quadrupedal robots, which require not only a robust locomotion controller itself but also consider the stability of such a controller over different terrains.
@@
The central contribution of this work is the development of a hierarchical RGBD-based traversability-aware navigation-locomotion framework using reinforcement learning. This framework enables the robot to learn to reach goals, avoid obstacles, and minimize trajectory traversability cost from color and depth information in the real world within 15 minutes. We leverage the value function obtained during the training of the low-level locomotion controller using the actor-critic method. To develop a planner that interprets terrain texture and geometry in the real world, we enhance an off-policy RL framework to learn from both online data and offline demonstrations. Validated through simulation and real-world experiments, our framework demonstrates better performance in terms of better estimation of the traversability, and improved generalization to new environments compared to baseline methods.
@@ II-A Traversability estimation
To estimate the terrain traversability cost, some previous work directly extracts features from the environment. These methods usually define traversability by empirically judging the cost from meshes [4], point clouds [5], or depth images [6]. Others try to classify the terrains from semantic categories [7, 8, 9], often mapping semantic features to traversability cost based on human-provided labels. These methods typically rely on human prior knowledge rather than the robot’s own experiences [10]. While these estimators are robot-agnostic and easily transferable across different embodiments, they fail to consider the robot’s real interaction experiences.
To consider the robot’s experience while traversing different terrains, previous work has suggested various metrics. The most straightforward approach is to consider control performance metrics such as tracking error of commands (e.g., velocity) [11] or IMU variance[12]. These quantitative metrics are used as labels for supervised learning of environment representations. Other approaches extend this research by learning estimators to predict traversal time [13], selected environment physics parameters [14], or collision and fall indicators [15] from a history of robot’s proprioception feedback. While existing methods link terrain traversability to specific metrics, they overlook the full-body performance of legged robots with a controller, such as stability and energy efficiency. We address this by developing a proprioception-based traversability estimator using the value function of the RL-based locomotion controller.
@@ II-B Learning based navigation
While numerous methods can tackle the planning problem for quadrupedal robots [16], we focus on learning-based methods that provide a planning policy directly from the robot’s vision input. Some prior work opts to use supervised learning by imitating expert trajectories or demonstrations provided by humans, such as [17]. In these methods, the robot is trained to reason expert actions from vision-based observations and mimic the expert’s strategies. However, the traversability cost can differ significantly across different embodiments. If the robot only learns from demonstrations provided by others (e.g., humans), it cannot update and optimize its planning strategy based on its own experiences: the planning policy is limited to offline data.
Reinforcement learning offers an approach that allows the robot to update its planning strategy based on its interaction with the environment, i.e., to learn from online data. Usually, RL requires extensive data through trial and error, therefore, some prior attempts leverage photo-realistic simulation to train the robot extensively and directly transfer it to the real world, primarily in indoor environments such as [18]. However, for outdoor environments where large sim-to-real gaps exist, due to factors like varying lighting conditions and terrain types [11], zero-shot transfer of a color-vision-based policy from sim to real is still challenging.
Therefore, some other attempts try to train the planning policy directly with data collected in the real world, such as on racing cars with states-based data [19] or vision-based data [10]. For more complex robots, like quadrupedal robots, there have been attempts to train state-based control policies from scratch [20]. However, using RL to deploy real-world training for quadrupedal robots to extract vision features, which are in a much higher dimension, remains an open question. This requires a sampling-efficient yet capable online learning algorithm to learn from high-dimension vision data. To tackle this and obtain a traversability-aware planning policy for real-world deployment, we enable the quadrupedal robot to take advantage of both online data (its own experience) and offline data, such as expert demonstrations and embeddings from a pretrained color image encoder.
@@ III-A Problem Formulation
  • Traversability Cost Minimization: The robot should be aware of the interplay between terrain traversability and locomotion capability, minimizing the traversability cost over trajectories. Terrain texture can be perceived by the robot’s color vision. The terrain’s traversability should depend on the robot’s own interaction experiences instead of predefined labels, semantic classifications, or human-centric experiences.
@@ III-B Framework
The proposed traversability-aware hierarchical-legged navigation framework consists of two parts: (1) a low-level locomotion controller $\pi^{c}$ for the quadrupedal robot, which can track varying velocity and turning commands over different terrains, and (2) a high-level planner $\pi^{p}_{\text{rgbd}}$ that considers the navigation goal, surroundings, and terrain textures to specify control commands. As illustrated in Fig. 2, the training of this framework consists of four stages. They are detailed below.
@@ III-B1 Stage 1: Training a Robust Locomotion Controller and a Traversability Estimator
Building the framework from scratch, we first obtain a robust locomotion controller $\pi^{c}$ for the quadrupedal robot using RL, as shown in Fig. 3. Our novel design includes the development of an estimator for overall control performance changes induced by changes in terrain traversability $T$ as detailed in Sec. IV-B.
@@ III-B2 Stage 2: Training a Depth-Based Navigation Planner as a Demonstrator
After obtaining the locomotion controller, we start training the planner $\pi^{p}_{\text{depth}}$ on top of it. We aim to train a zero-shot deployable planner with only depth input in simulation, which will later act as a demonstrator during real-world data collection. The objective of this planner is to drive the robot to a target location while avoiding nearby obstacles, which are perceived through depth images. Details will be introduced in Sec. V-A.
@@ III-B3 Stage 3: Data Collection in the Real World
To increase sample efficiency for the real-world training, we first collect rollouts in the real world with the trained depth-based planner $\pi^{p}_{\text{depth}}$ and a human expert. During these rollouts, data is collected in the form of transition pairs: $\langle\mathbf{o}^{p}_{\text{rgbd},t},\mathbf{a}^{p}_{t},r^{p}_{t},\mathbf{o}^{p}_{\text{rgbd},t+1},\mathbf{a}^{p}_{t+1}\rangle$. These rollouts allow us to collect extended observations $\mathbf{o}^{p}_{\text{rgbd}}$, including RGB information that provides additional visual texture features of terrains and surroundings. The reward $r^{p}_{t}$ collected during rollouts includes the estimated traversability cost $T$ based on the robot’s real experience interacting with the terrains. The details of this stage are discussed in Sec. V-B.
@@ IV-A Legged Locomotion Control over Complex Terrains
We first leverage RL to develop a robust blind locomotion controller for the quadrupedal robot, enabling it to traverse different terrains while tracking varying velocities. We use an actor-critic algorithm, PPO [23], where the goal-conditioned actor policy $\pi^{c}$ and critic network $V^{c}$ are trained jointly in Isaac Gym [24] with randomized robot dynamics and terrain features. We adopt a symmetric actor critc structure to ensure that the critic can also be inferenced on the real-world robot. The controller is designed to track varying goal commands $\mathbf{g}_{k}$, including walking velocities in the sagittal and lateral directions $\dot{q}^{d}_{[x,y],k}$ and turning rate $\dot{q}^{d}_{\psi,k}$.
IV-B Traversability Estimation Based on Value Function
Our design is based on the intuition that if the robot traverses challenging terrains easily, the estimated traversability cost will remain low, meaning the traversability should be scaled according to the locomotion controller’s strength. As a result, the critic network’s estimated value, representing expected future rewards, is inversely related to the terrain’s traversability cost.
A key issue with using the critic network directly is its sensitivity to goal command changes $\mathbf{g}_{k}$, causing value significant changes even on flat ground. This problem is further compounded by the sim-to-real gap which produces inaccurate predictions and potential overfitting to the simulation data.
@@ IV-B Traversability Estimation Based on Value Function
As shown in Fig. 3, it contains three steps to obtain the final estimated traversability cost $T$ to use in the real world, as detailed in the following.
@@
Training Uncertainty-Aware Value Estimator: As shown in Fig. 3 (iii), we use the parameters of the pre-trained $V^{c}$ and add a dropout layer right before its output layer, initializing a value estimation network $V_{\text{terrain}}^{c}$. For inference, we perform $n$ forward passes with the same observation ($\mathbf{o}_{k}$) and condition ($\mathbf{g}_{k}$). We then calculate the mean of the predictions $\text{Mean}(V_{\text{terrain}}^{c,(i=1:n)}(\mathbf{o}_{k}|\mathbf{g}_{k}))$ as the estimated return, and the standard deviation $\text{Var}(V_{\text{terrain}}^{c,(i=1:n)}(\mathbf{o}_{k}|\mathbf{g}_{k}))$ as the uncertainty metric, inspired by [25]. This procedure acts like ensemble methods, where disagreement among multiple predictions indicates higher uncertainty, suggesting that the predicted return is less trustworthy and should be treated with caution.
Figure 3: The training framework for locomotion and traversability estimator. (i) The goal-conditioned controller (actor policy $\pi^{c}$) is trained to perform quadrupedal trotting gaits while tracking velocity commands $\mathbf{g}^{c}_{k}$ over various terrains. (ii) A baseline value function $V^{c}_{\text{flat}}$ is trained to estimate the return on flat grounds. (iii) We add a dropout layer before the output of $V^{c}$ and fine-tune the critic on complex terrains to create an uncertainty-aware estimator ($V^{c}_{\text{terrain}}$). (iv) In real-world deployment, we perform $n$ inferences using $V^{c}_{\text{terrain}}$ for the same observation $\mathbf{o}^{c}_{k}$ and treat the Variance $\sigma$ of these inferences measures as uncertainty. The mean value is adjusted by subtracting the baseline $V^{c}_{\text{flat}}$ to reduce bias induced by the given goal $\mathbf{g}^{c}_{k}$. The negative of this difference is then multiplied by an adaptive weight $(a-b\sigma)$, resulting in the traversability estimation $T_{\text{value}}$.
@@
This method of evaluating the difference between a baseline return estimation and an estimator for changes in the environment is not limited to traversability estimation as focused in this work. It can serve as a general evaluator of RL-based control performance. For example, it can also assess the sim-to-real gap: by comparing a baseline estimator for nominal dynamics parameters with an estimator for different dynamics parameters under the same given goal.

### v2

@@ Abstract
The enhanced mobility brought by legged locomotion empowers quadrupedal robots to navigate through complex and unstructured environments. However, optimizing agile locomotion while accounting for the varying energy costs of traversing different terrains remains an open challenge. Most previous work focuses on planning trajectories with traversability cost estimation based on human-labeled environmental features. However, this human-centric approach is insufficient because it does not account for the varying capabilities of the robot locomotion controllers over challenging terrains. To address this, we develop a novel traversability estimator in a robot-centric manner, based on the value function of the robot’s locomotion controller. This estimator is integrated into a new learning-based RGBD navigation framework. The framework employs multiple training stages to develop a planner that guides the robot in avoiding obstacles and hard-to-traverse terrains while reaching its goals. The training of the navigation planner is directly performed in the real world using a sample efficient reinforcement learning method that utilizes both online data and offline datasets. Through extensive benchmarking, we demonstrate that the proposed framework achieves the best performance in accurate traversability cost estimation and efficient learning from multi-modal data (including the robot’s color and depth vision, as well as proprioceptive feedback) for real-world training. Using the proposed method, a quadrupedal robot learns to perform traversability-aware navigation through trial and error in various real-world environments with challenging terrains that are difficult to classify using depth vision alone. Moreover, the robot demonstrates the ability to generalize the learned navigation skills to unseen scenarios. Video can be found at https://youtu.be/RSqnIWZ1qks.
Index Terms:
Legged robot navigation, traversability estimation, reinforcement learning, real-world training
@@ I Introduction
Figure 1: Our proposed framework enables a quadrupedal robot to learn to avoid hard-to-traverse terrains (such as muddy areas covered by leaves) and obstacles from its own interactions through training directly in the real world. The robot learns to utilize its onboard color and depth vision sensors to identify challenging terrains and obstacles while navigating toward the goals, marked by the blue dashed circles. The inclusion of RGB images allows the robot to identify additional terrain textures that are difficult to perceive with depth images alone during the navigation.
@@
However, estimating the traversability cost of terrain and planning to avoid high-cost regions for legged robots is a challenging problem. It requires the robot to interpret visual textures and corresponding physical properties, such as friction and elasticity, of different terrains. For example, when humans encounter a muddy road with high traversability costs, we first see the mud’s texture and infer its difficulty based on previous experience. We highlight the importance of learning from previous experience as it’s difficult to accurately estimate the terrain’s actual cost until we traverse it ourselves. For legged robots to develop similar intelligence, they must address these multi-modal learning challenges, integrating data from various sources. To estimate terrain traversability cost, the robot needs to fuse color data (terrain texture), depth data (identifying non-traversable regions like obstacles), and proprioceptive feedback (implicitly encoding the physics) in real time and the real world. This problem is further complicated by the complex dynamics of quadrupedal robots, which require not only a robust locomotion controller itself but also consider the stability of such a controller over different terrains.
@@
The central contribution of this work is the development of a hierarchical RGBD-based traversability-aware navigation-locomotion framework using reinforcement learning. This framework enables the robot to learn to reach goals, avoid obstacles, and minimize trajectory traversability cost from color and depth information in the real world within 15 minutes. To quantitatively evaluate terrain traversability cost, we leverage the value function obtained during the training of the low-level locomotion controller using the actor-critic method. Traditionally, the critic is only used during training and not during deployment. However, in this work, we deploy the trained value function to evaluate control performance across different terrains, thus representing the terrain traversability cost. This estimator, based solely on the robot’s proprioceptive feedback, can also be used in different modules. To develop a planner that interprets terrain texture and geometry in the real world, we enhance an off-policy RL framework to learn from both online data and offline demonstrations. Validated through simulation and real-world experiments, our framework demonstrates better performance in terms of better estimation of the traversability, and improved generalization to new environments compared to baseline methods.
@@ II-A Traversability estimation
To estimate the terrain traversability cost, some previous work directly extracts features from the environment. These methods usually define traversability by empirically judging the cost from meshes [4], point clouds [5], or depth images [6]. Others try to classify the terrains from semantic categories [7, 8, 9], often mapping semantic features to traversability cost based on human-provided labels. These methods typically rely on human prior knowledge rather than the robot’s own experiences [10]. While these estimators are robot-agnostic and easily transferable across different embodiments, they fail to consider the robot’s real interaction experiences. For example, some terrains may be difficult for lightweight robots but manageable for heavier ones, and vice versa.
To consider the robot’s experience while traversing different terrains, previous work has suggested various metrics. The most straightforward approach is to consider control performance metrics such as tracking error of commands (e.g., velocity) [11] or IMU variance[12]. These quantitative metrics are used as labels for supervised learning of environment representations. Other approaches extend this research by learning estimators to predict traversal time [13], selected environment physics parameters [14], or collision and fall indicators [15] from a history of robot’s proprioception feedback. While these methods link terrain traversability to specific control performance metrics, they do not capture the whole-body performance (including velocity tracking, base stability, motion smoothness, energy efficiency, and more) of a legged robot’s full-order dynamics driven by its controller. In this work, we address this by developing a traversability estimator from proprioception based on the value function of the RL-based locomotion controller.
@@ II-B Learning based navigation
While numerous methods can tackle the planning problem for quadrupedal robots [16], we focus on learning-based methods that provide a planning policy directly from the robot’s vision input. Some prior work opts to use supervised learning by imitating expert trajectories or demonstrations provided by humans, such as [17, 18, 19]. In these methods, the robot is trained to reason expert actions from vision-based observations along demonstrated trajectories. This approach enables the robot to acquire behaviors quickly by mimicking successful strategies. However, the traversability cost can differ significantly across different embodiments. If the robot only learns from demonstrations provided by others (e.g., humans), it cannot update and optimize its planning strategy based on its own experiences: the planning policy is limited to offline data.
Reinforcement learning offers an approach that allows the robot to update its planning strategy based on its interaction with the environment, i.e., to learn from online data. Usually, RL requires extensive data through trial and error, therefore, some prior attempts leverage photo-realistic simulation to train the robot extensively and directly transfer it to the real world, primarily in indoor environments such as [20, 21, 22, 23]. However, for outdoor environments where large sim-to-real gaps exist, due to factors like varying lighting conditions and terrain types [11], zero-shot transfer of a color-vision-based policy from sim to real is still challenging.
Therefore, some other attempts try to train the planning policy directly with data collected in the real world, such as on racing cars with states-based data [24] or vision-based data [10]. For more complex robots, like quadrupedal robots, there have been attempts to train state-based control policies from scratch [25, 26]. However, using RL to deploy real-world training for quadrupedal robots to extract vision features, which are in a much higher dimension, remains an open question. This requires a sampling-efficient yet capable online learning algorithm to learn from high-dimension vision data. To tackle this and obtain a traversability-aware planning policy for real-world deployment, we enable the quadrupedal robot to take advantage of both online data (its own experience) and offline data, such as expert demonstrations and embeddings from a pretrained color image encoder.
@@ III-A Problem Formulation
  • Traversability Cost Minimization: The robot should be aware of the interplay between terrain traversability and locomotion capability, minimizing the traversability cost over trajectories. The terrain’s traversability should depend on the robot’s own interaction experiences instead of predefined labels, semantic classifications, or human-centric experiences.
@@ III-B Framework
The proposed traversability-aware hierarchical-legged navigation framework consists of two parts: (1) a low-level locomotion controller $\pi^{c}$ for the quadrupedal robot, which can track varying velocity and turning commands over different terrains, and (2) a high-level planner $\pi^{p}_{\text{rgbd}}$ that considers the navigation goal, surroundings, and terrain textures to specify control commands. As illustrated in Fig. 2, the training of this framework consists of four stages.
@@ III-B1 Stage 1: Training a Robust Locomotion Controller and a Traversability Estimator
Building the framework from scratch, we first obtain a robust locomotion controller $\pi^{c}$ for the quadrupedal robot using RL, as shown in Fig. 3. Our novel design includes the development of an estimator for overall control performance changes induced by changes in terrain traversability $T$. This estimator is based on the critic network $V^{c}$ obtained during the training of the control policy. Several design choices are implemented to ensure this estimator’s functionality for traversability estimation on the real robot, detailed in Sec. IV-B. Both the controller and estimator are trained in simulation.
@@ III-B2 Stage 2: Training a Depth-Based Navigation Planner as a Demonstrator
After obtaining the locomotion controller, we start training the planner $\pi^{p}_{\text{depth}}$ on top of it. We aim to train a zero-shot deployable planner with only depth input in simulation, which will later act as a demonstrator during real-world data collection. The objective of this planner is to drive the robot to a target location while avoiding nearby obstacles, which are perceived through depth images. This planner will act as a demonstrator for autonomous data collection in the real world. Details will be introduced in Sec. V-A.
@@ III-B3 Stage 3: Data Collection in the Real World
To increase sample efficiency for the real-world training, we first collect rollouts in the real world with the trained depth-based planner $\pi^{p}_{\text{depth}}$. Additionally, our framework allows for the inclusion of human demonstrations, where human experts directly navigate the robot. During these rollouts, data is collected in the form of transition pairs: $\langle\mathbf{o}^{p}_{\text{rgbd},t},\mathbf{a}^{p}_{t},r^{p}_{t},\mathbf{o}^{p}_{\text{rgbd},t+1},\mathbf{a}^{p}_{t+1}\rangle$. These rollouts allow us to collect extended observations $\mathbf{o}^{p}_{\text{rgbd}}$, including RGB information that provides additional visual texture features of terrains and surroundings. The reward $r^{p}_{t}$ collected during rollouts includes the estimated traversability cost $T$ based on the robot’s real experience interacting with the terrains. The details of this stage are discussed in Sec. V-B.
@@ IV-A Legged Locomotion Control over Complex Terrains
We first leverage RL to develop a robust blind locomotion controller for the quadrupedal robot, enabling it to traverse different terrains while tracking varying velocities. We use an actor-critic algorithm, PPO [29], where the goal-conditioned actor policy $\pi^{c}$ and critic network $V^{c}$ are trained jointly in simulation. We adopt a symmetric actor critc structure to ensure that the critic can also be inferenced on the real-world robot. The controller is designed to track varying goal commands $\mathbf{g}_{k}$, including walking velocities in the sagittal and lateral directions $\dot{q}^{d}_{[x,y],k}$ and turning rate $\dot{q}^{d}_{\psi,k}$. The reward is designed to accomplish different control sub-objectives with different weights, including tracking given commands, following parameterized reference motion obtained by a central pattern generator [30], and several auxiliary terms including stabilizing base orientation, minimizing base vertical velocity, joint velocity, and acceleration. We randomize robot dynamics and terrain features and the training is done in Isaac Gym [31].
IV-B Estimation of Control Performance Based on Value Function
Our design originates from the following intuition: if the robot can traverse challenging terrains easily, the estimated traversability cost will still be low. This indicates that the level of traversability should be scaled according to the strength of the locomotion controller. The critic network estimates the value, Interestingly, the traversability cost of the terrain can be inversely related to the estimated value from the critic network of locomotion representing the expected future accumulated reward.
However, directly using the critic network for real-world evaluation of traversability is challenging. A frequent problem we encountered is that the estimated value is also sensitive to changes in the given commands $\mathbf{g}_{k}$ using the goal-conditioned controller. For example, the critic return varies even when the robot locomotes on flat ground due to imperfect velocity tracking given different commands $\mathbf{g}_{k}$. This issue is further compounded by the sim-to-real gap: the critic may produce incorrect value predictions due to domain shifts, potentially overfitting to the simulation.
@@ IV-B Estimation of Control Performance Based on Value Function
As explained in Algorithm 1, it contains three steps to obtain the final estimated traversability cost $T$ to use in the real world, as detailed in the following.
@@
Training Uncertainty-Aware Value Estimator: As shown in Fig. 3 (iii), we use the parameters of the pre-trained $V^{c}$ and add a dropout layer right before its output layer, initializing a value estimation network $V_{\text{terrain}}^{c}$.
For inference, we perform $n$ forward passes with the same observation ($\mathbf{o}_{k}$) and condition ($\mathbf{g}_{k}$). We then calculate the mean of the predictions $\frac{1}{n}\sum_{i=1}^{n}V_{\text{terrain}}^{c,(i)}(\mathbf{o}_{k}|\mathbf{g}_{k})$ as the estimated return, and the standard deviation $\text{Var}(V_{\text{terrain}}^{c,(i=1:n)}(\mathbf{o}_{k}|\mathbf{g}_{k}))$ as the uncertainty metric, inspired by [32]. This procedure acts like ensemble methods, where disagreement among multiple predictions indicates higher uncertainty, suggesting that the predicted return is less trustworthy and should be treated with caution.
Remark 1
The dropout layer is only added after the actor and critic networks are trained. The reason we do not concurrently train the critic with dropout is that the added dropout layer introduces a higher variance and fitting error, making the training for the actor network unstable.
Figure 3: The training framework for locomotion and traversability estimator. (i) The goal-conditioned controller (actor policy $\pi^{c}$) is trained to perform quadrupedal trotting gaits while tracking velocity commands $\mathbf{g}^{c}_{k}$ over various terrains. (ii) After obtaining $\pi^{c}$, a baseline value function $V^{c}_{\text{flat}}$ is trained to estimate the return while the robot is walking on flat ground using the same actor policy. (iii) To create an uncertainty-aware estimator ($V^{c}_{\text{terrain}}$), we add a dropout layer before the output of $V^{c}$ and fine-tune it on complex terrains. (iv) In real-world deployment, we perform $n$ inferences using $V^{c}_{\text{terrain}}$ for the same goal $\mathbf{g}_{k}$ and observation $\mathbf{o}^{c}_{k}$. The standard deviation $\sigma$ of these inferences measures uncertainty. The mean value is adjusted by subtracting the baseline $V^{c}_{\text{flat}}$ to reduce bias induced by the given goal $\mathbf{g}^{c}_{k}$. The negative of this difference is then multiplied by an adaptive weight $(a-b\sigma)$, where $a$ and $b$ are tunable parameters, resulting in the traversability estimation $T_{\text{value}}$.
Algorithm 1 Traversability Estimation
  GIVEN Trained locomotion controller and its critic network $\pi^{c}$, $V^{c}$
  Output $T_{\text{value}}$ with uncertainty
  // Train baseline value estimator $V^{c}_{\text{flat}}$
[…]
            Sample $\mathbf{g}^{c}_{k}=\dot{q}^{d}_{[x,y,\psi],k}$ under flat ground
            Observe $\mathbf{o}^{c}_{k}$
            Run $\pi^{\text{c}}(\mathbf{o}^{c}_{k}|\mathbf{g}^{c}_{k})$ and collect future return $R_{k}$
        end for
        Update $V^{c}_{\text{flat}}(\mathbf{g}^{c}_{k})$ to fit return $R_{k}$
  end for
  // Train uncertain-aware estimator $V^{c}_{\text{terrain}}$
  Initialize Add a dropout layer next to the final layer of $V^{c}$ and obtain initial $V^{c}_{\text{terrain}}$
[…]
            Sample $\mathbf{g}^{c}_{k}=\dot{q}^{d}_{[x,y,\psi],k}$ and complex terrain types
            Observe $\mathbf{o}^{c}_{k}$
            Run $\pi^{c}(\mathbf{o}^{c}_{k}|\mathbf{g}^{c}_{k})$ and collect future return $R_{k}$
        end for
        Update $V^{c}_{\text{terrain}}(\mathbf{o}^{c}_{k}|\mathbf{g}^{c}_{k})$ with dropout to fit return $R_{k}$
  end for
  // Estimate $T_{\text{value}}$ with uncertainty
  Observe $\mathbf{o}^{c}_{k}$, $\mathbf{g}^{c}_{k}$
  for $i=1:n$ inference
        Run $i$-th forward pass on $V^{c,(i)}_{\text{terrain}}(\mathbf{o}^{c}_{k}|\mathbf{g}^{c}_{k})$ with dropout
  $T_{\text{value}}\leftarrow$ $V^{c}_{\text{flat}}(\mathbf{g}^{c}_{k})-\text{mean}(V^{c,(i=1:n)}_{\text{terrain}}(\mathbf{o}^{c}_{k}|\mathbf{g}^{c}_{k}))$
  with uncertainty $\sigma(V^{c,(i=1:n)}_{\text{terrain}}(\mathbf{o}^{c}_{k}|\mathbf{g}^{c}_{k}))$
  end for
@@
This method of evaluating the difference between a baseline return estimation and an estimator for changes in the environment is not limited to traversability estimation as focused in this work. It can serve as a general evaluator of RL-based control performance. For example, it can also assess the sim-to-real gap: by comparing a baseline estimator for nominal dynamics parameters with an estimator for different dynamics parameters under the same given goal, a significant gap indicates a large sim-to-real discrepancy.